OpenStudy (anonymous):

convert prefix 20 ms^-1 ---> ? kmh^-1

OpenStudy (anonymous):

1kn=1000m 1h=3600s hence 1m=1/1000km 1s=1/3600s so 20m/s=20 * 3600/1000=20*18/5=72kph

OpenStudy (anonymous):

why 20*3600?

OpenStudy (anonymous):

There's 60 seconds in a minute, and 60 minutes in an hour. 60*60 = 3600. So 20 metres per second = 20*60*60 = 20*3600 metres per hour. Divide by 1000 to convert to kilometres. An easier conversion from m/s -> km/h is simply multiply by 3.6, I'm sure you can see the shortcut from the above.